find the solutions of each system
3x+2y=5
4x+3y=5

Answer by fractalier(6550)

Let us multiply the top one by 4 and the bottom one by 3 and subtract them...we get
3x + 2y = 5
4x + 3y = 5
which becomes
12x + 8y = 20
-(12x + 9y = 15)
and thus
-y = 5
and
y = -5
so that when we substitute this into the first equation we get
3x + 2(-5) = 5
3x - 10 = 5
3x = 15
x = 5
Our solution is (5, -5)